OpenStudy (anonymous):

Help me please !!!!! I have f(x) = 6x - x^2 - 6sqrt(a+b) determine: (f(sqrt(a+b) - f(1)) / sqrt(a+b) = ? Thank you for your help

myininaya (myininaya):

\[f(\sqrt{a+b})=6\sqrt{a+b}-(\sqrt{a+b})^2-6\sqrt{a+b} \]

myininaya (myininaya):

simplify this

OpenStudy (anonymous):

\[f(x)=6x-x ^{2}-6\sqrt{a+b} \]

OpenStudy (anonymous):

but i need to find: \[\frac{ f(\sqrt{a+b})-f(1) }{ \sqrt{a+b}}=?\]

myininaya (myininaya):

Yes I'm asking you to find that one thing first.

myininaya (myininaya):

Just simplify what I wrote above and we are going to plug that into your expression that you are trying to evaluate

myininaya (myininaya):

 This is our main goal to evaluate But right now I'm having you work on the little parts of it first like this circled thing: